OpenStudy (anonymous):

f(x) = 3x + 2; g(x) = 3x - 5 find f/g

OpenStudy (anonymous):

please help

hero (hero):

\[\frac{f(x)}{g(x)} = \frac{3x + 2}{3x - 5} = \frac{3x -5 + 7}{3x -5} = \frac{3x - 5}{3x - 5} + \frac{7}{3x - 5} = 1 + \frac{7}{3x - 5}\]

OpenStudy (anonymous):

Hi Hero

hero (hero):

Hi @Marlins0412

OpenStudy (anonymous):

so the answer to this problem is 1-7/3x-5?

hero (hero):

Now just think about it for a minute and realize for yourself what the correct answer is.

hero (hero):

Also realize that no one even bothered to dispute my result

OpenStudy (anonymous):

Ok but what would be the domain though?

hero (hero):

Any x as long as x ≠ 5/3

OpenStudy (anonymous):

can you help with 1 more please ?

hero (hero):

Okay

OpenStudy (anonymous):

f(x) = 4x + 7, g(x) = 3x^2 Find(fg)(x)

hero (hero):

\[(fg)(x) = f(x) \dot\ g(x) = (4x + 7)\dot\ (3x^2) = 4x(3x^2) + 7(3x^2) = 12x^3 + 21x^2\]

OpenStudy (anonymous):

Thanks so much if you don't mind i have a few more but its totally up to you!

hero (hero):

I'm more interested in understanding whether or not you're getting this.

OpenStudy (anonymous):

As you do these instructions i am completely understanding your so good! May you help with more?

hero (hero):

I guess, but like I said, I'm more interested in what you can do after I've shown you.

OpenStudy (anonymous):

I understand what directions you give. THe next problem is Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y=(7/x^2)+10

hero (hero):

f(x) = x + 10 g(x) = 7/x^2 f(g(x)) = f(7/x^2) = (7/x^2) + 10

OpenStudy (anonymous):

ok so you did the square of each one i see what you did. Wow your a teacher i bet!!

OpenStudy (anonymous):

THe next one i need help with is f(x) = 3x + 9, g(x) = 5x^2 Find (f+g)(x)

hero (hero):

How do you know I'm not a HS Student? When I was in high school, I did these exactly the same.

OpenStudy (anonymous):

wow you must be a brilliant person!

hero (hero):

Add you have to do is add f(x) + g(x) = (3x + 9) + (5x^2)

hero (hero):

Add them and let me know what you get.

OpenStudy (anonymous):

well im thinking the answer is 15x^3+45x

OpenStudy (anonymous):

5x^2 + 3x + 9 is what i got when adding...

hero (hero):

Yes, correct. You can only combine like terms. Since there are no like terms to combine, you simply express the addition as a quadratic expression.

hero (hero):

So the answer is 5x^2 + 3x + 9 (not 15x^3 + 45x)

OpenStudy (anonymous):

right so the first answer i got was multiplying//

OpenStudy (anonymous):

do you know how to do inverses?

hero (hero):

"Do I know how to do inverses?"...You're funny

hero (hero):

The question is, "Do you know how to do them?".

OpenStudy (anonymous):

Find the inverse of the function. f(x)=x^3-3

OpenStudy (anonymous):

Im saying for you to help me out with them...

hero (hero):

`1.` Replace f(x) with y `2.` Add three to both sides `3.` Cube root both sides `4.` Swap x and y `5.` Replace y with \(f^{-1}{(x)}\)

OpenStudy (anonymous):

x+2/7?

OpenStudy (anonymous):

are you still there?

hero (hero):

If that's what you got, you didn't follow my steps exactly.

OpenStudy (anonymous):

is it x-2/7 ?

hero (hero):

Of course it isn't

OpenStudy (anonymous):

i know its f^_1x but after that is where i need the help

hero (hero):

Bro, just follow the steps I gave you.